OpenStudy (anonymous):

Consider the basis B of R2 consisting of vectors 5 -6 and -2 -2 Find x in R2 whose coordinate vector relative to the basis B is [x]B = [-6] [ 2 ] and please explain to me how you found it. Thank You

OpenStudy (anonymous):

The idea is to find a linear combination a_1(5, -6) + a_2(-2, -2) = (-6, 2) It boils down to a system of equations: Take the augmented matrix: \[ \left[ \begin {array}{ccc} 5&-2&-6\\ -6&-2&2 \end {array} \right] \] Reduced form: \[ \left[ \begin {array}{ccc} 1&0&-{\frac {8}{11}}\\ 0 &1&{\frac {13}{11}}\end {array} \right] \] -(8/11)*(5, -6) + (13/11)*(-2, -2) = (-6, 2)

OpenStudy (anonymous):

the answer is supposed to be another 2 by 1 matrix

OpenStudy (anonymous):

\[\left(\begin{matrix}-\frac{8}{11} \\ \frac{13}{11}\end{matrix}\right)\]

OpenStudy (anonymous):

Let me try to rewrite it to show you. You are solving a system of equations.

OpenStudy (anonymous):

The solution is the column vector (2 by 1 matrix) \[ \left[ \begin {array}{c} -{\frac {8}{11}}\\ {\frac {13}{11}}\end {array} \right] \]

OpenStudy (anonymous):

\[5x_1 -2x_2 = -6\]\[-6x_1 - 2x_2 = 2\]

OpenStudy (anonymous):

Solve for x_1 and x_2 to find the coordinates of the coordinate vector.